At a height of 312 km above the Earth's surface, an astronaut finds that the atmospheric pressure is about 1.35E-8 mmHg and the temperature is 585 K. How many molecules of gas are there per milliliter at this altitude?
1
Expert's answer
2012-10-08T10:25:39-0400
pV=nRT
n/V = p/RT = 1.35E-8 mmHg / (62364 mmHg*mL/mol*K * 585 K) = 3.7E-16
mol/mL
N/V = n/V * N(A) = 3.7E-16 * 6.022E23 = 2.23E8 molecules/mL =
222'835'659 molecules/mL